# Towards Open-World Gesture Recognition

### v1

@@ Abstract
Static machine learning methods in gesture recognition assume that training and test data come from the same underlying distribution. However, in real-world applications involving gesture recognition on wrist-worn devices, data distribution may change over time. We formulate this problem of adapting recognition models to new tasks, where new data patterns emerge, as open-world gesture recognition (OWGR). We propose leveraging continual learning to make machine learning models adaptive to new tasks without degrading performance on previously learned tasks. However, the exploration of parameters for questions around when and how to train and deploy recognition models requires time-consuming user studies and is sometimes impractical. To address this challenge, we propose a design engineering approach that enables offline analysis on a collected large-scale dataset with various parameters and compares different continual learning methods. Finally, design guidelines are provided to enhance the development of an open-world wrist-worn gesture recognition process.
@@ Introduction
Wrist-worn gesture recognition using inertial measurement unit (IMU) signals offers a convenient, always-on interface for various applications in mixed reality [21, 52, 40, 44]. Currently, most hand gesture recognition algorithms are optimized for “closed-world” settings, where training and test data come from the same underlying distribution [51]. However, in real-world applications [3, 41], new gesture data continuously arrives with changing characteristics, gesture data may change over time, and entirely new data patterns can emerge. We identify this problem setting as open-world gesture recognition (OWGR). Static machine learning, which trains and deploys a recognition model only once, cannot effectively tackle this problem. What we seek is an adaptive machine learning method that can continuously train and deploy the model on newly emerging data, as illustrated in Figure Towards Open-World Gesture Recognition.
Various approaches can be employed for adaptive machine learning. Simply retraining a recognition model with the entire joint dataset (past data and new data from new tasks) is sometimes infeasible due to limited computational power on embedded devices. Storing the entire past dataset is also challenging due to the limited memory of the device and privacy concerns. On the other hand, if a model trained from past data is naively finetuned on a new task, that model will dramatically decrease the recognition performance on the old tasks [45]. This is called catastrophic forgetting. Continual learning (also called lifelong learning) methods are specifically designed to alleviate catastrophic forgetting by balancing the trade-off between plasticity (transfer knowledge from old task to new task) and stability (catastrophic forgetting).
We identify three real-world cases that fall under OWGR (shown in Figure Towards Open-World Gesture Recognition): The first one is new context, where the gesture recognition model adaptively learns to recognize existing gestures under new contexts, such as performing pinch when walking and running, without forgetting old contexts such as performing pinch when standing. The second case is new gesture when the recognition model must learn to recognize a new gesture. The third case is new user, where the recognition model learns to recognize existing gestures as performed by new users, such as when a device is shared between users, or a newly-bought device calibrates to a user for the first time. In all cases, we wish to preserve performance on previously learned tasks (to avoid catastrophic forgetting) while also learning the new task. Here, a new task may refer to a new context, a new gesture, or a new user. In Section 3, we present a detailed formulation of the open-world gesture recognition problem and its associated real-world scenarios. This stands as the paper’s primary contribution since we are the first in defining this problem and methodically pinpointing the cases.
@@
In static machine learning, we only need to log the data from the device, optimize the recognition model on the logged data to best fit the data once, and then deploy a trained model onto the device. However, in an adaptive machine learning process for OWGR, we additionally must employ several other stages as illustrated in Figure 1. This OWGR process includes three additional critical stages, which require careful design considerations: (1) when to update the model, and (2) how to update the model. When a developer of a wrist-worn open-world gesture recognition system seeks to assess and refine their approach, particularly when delving into varied use cases like the three we previously mentioned, challenges emerge. These stem from the intricacies of conducting actual user studies, given the numerous parameters. For instance, evaluating an OWGR process under new context via a user study could necessitate a few participants performing gestures across many distinct contexts in the real world instead of in the lab. To optimize a functional OWGR process, multiple such evaluations might be needed, making it extremely difficult to conduct and validate.
@@
Following the practice of Kristensson et al. [25], which proposes using a design engineering approach to assess a process that is challenging to evaluate via user studies, we have adopted a similar design engineering method. This strategy outlines a functional design of open-world gesture recognition and examines its conceptual design in which the key function of the process (Updating Policy) is translated into six potential function carriers (i.e. five different implementations of continual learning methods and one finetuning baseline). We then distinguish between controllable and uncontrollable parameters and assess the process’s effectiveness through quantitative envelope analysis. Such a method lets us assess the parameters offline, thus sidestepping the need for on-ground deployments and tests with real users. To this end, we first collected a large-scale dataset using the inertial measurement unit (IMU) sensor from wrist-worn devices. To capture diversity in the aforementioned three real-world cases, the dataset contains 50 users performing four different gestures in 25 different contexts. We then construct a surrogate data model and a surrogate task model, identify their controllable and uncontrollable parameters, and subsequently perform an envelope analysis by varying these parameters. This illustrates that using a continual learning method for open-world gesture recognition significantly improves gesture recognition accuracy and reduces catastrophic forgetting compared to a finetuning method. This forms our second contribution to the paper.
Despite employing a design engineering methodology to assess and refine the OWGR process, designers are often restricted by the limited access to collect large-scale data and the limited availability of computational resources. This is because envelope analysis remains computationally demanding due to its intensive simulation nature, necessitating parameter sweeping. In this paper, the collection of this extensive dataset and our subsequent offline examinations, which encompassed the training and evaluation of different continual learning methods, imposed a significant resource burden, consuming over 40,000 GPU hours. Consequently, designers with constrained resources find it challenging to optimize the OWGR process. Our objective, therefore, is to distill essential design guidelines based on our engineering methodology’s outcomes to guide designers working on OWGR. The dataset we assembled holds generalizability due to its substantial user and context diversity. Additionally, as we employed representative continual learning methods, the design guidelines derived from this paper are widely applicable. Hence, this paper’s third contribution is the provision of distilled design guidelines for developing OWGR processes, underpinned by our design engineering approach using envelope analysis.
@@ 1.1 Wrist-Worn Devices with Mixed Reality
To overcome this limitation, wrist-worn devices have been employed for gestural interaction with mixed reality devices [2, 12]. Gesture recognition is conducted using these wrist-worn devices, eliminating the need for HMD cameras to detect hands and gestures. Using wrist-worn devices facilitates gestural interactions even when hands are under the table or in a pocket, offering a seamless user experience. Moreover, wrist-worn devices can recognize gestures in a broader variety of ways compared to front-facing cameras, which primarily rely on computer vision techniques. There has been extensive research exploring hand gesture recognition using a wide range of wrist-worn sensing modalities, such as RGB cameras [61, 17, 62], infrared (IR) ranging [33], intertial measurement unit (IMU) [27, 23, 16], acoustics [36, 26], electromyography (EMG) [45, 6, 34], electrical impedance tomography [68], pressure [11, 22], radar [29], stretch sensors [53], magnetic sensors [9, 37], and bio-capacitive effects [54]. IMU sensors are widely used for gesture recognition owing to their low cost and low power characteristics [20, 70, 63].
1.2 Wrist-Worn Gesture Recognition
@@
As for gesture recognition techniques, early trajectory-based gesture recognition methods (e.g., dynamic time warping (DTW) [30] and hidden Markov models (HMM) [35]) can recognize simple gesture trajectories such as drawing a shape  [35]). However, these methods do not work well for more complex and fine-grained gestures such as pinching or making a fist. Recently, researchers and practitioners mainly use data-driven approaches by collecting a labeled gesture dataset. They then either train traditional models such as SVM, trees, (e.g., [15, 18]) or deep learning models when the dataset is large enough (e.g., [17, 65]). However, most prior work focuses on recognizing a pre-defined gesture set. There are very few works addressing open-world gesture recognition. Xu et al. proposed a few-shot learning framework for gesture customization [63]. Shen et al. proposed a deep-learning model that can learn a new gesture with a synthetic dataset that is generated from a few data samples with a deep generative model [49, 47, 48]. While these approaches are useful in the new gesture case, they are not appropriate for the other two cases we identified, new context and new user, and they fail to address the catastrophic forgetting problem when old gestures are revisited.
Wang et al.  [59] identify a continual learning application for lifelong egocentric gesture recognition such that a VR system allows users to customize gestures incrementally. Our work differs in several ways. First, they only applied one continual learning method, iCaRl [39]. iCaRl requires large amounts of memory, which can be infeasible on wrist-worn devices. In contrast, we apply five different continual learning methods with varying trade-offs in our work. Second, they only applied continual learning to the new gesture scenario, whereas we also examine new context and new user. Third, they focused on egocentric gesture recognition with RGB and depth images as inputs, whereas we focus on wrist-worn gesture recognition with inputs being time-series IMU data. Lastly, we provide a systematic evaluation on various continual learning methods for different application scenarios, aiming to provide general design guidelines.
1.3 Design Engineering and Envelope Analysis
Design engineering offers a holistic method used in the creation and evolution of products and systems. This approach transitions from the initial stage of problem identification to tackling design-related aspects during the life cycle of a product or system, which includes its production, upkeep, and eventual decommissioning.
@@ 1.3 Design Engineering and Envelope Analysis
Validating certain systems, like a context-aware sentence retrieval system for AAC users, is extremely challenging. These systems require tailored setups and extended use before benefits emerge. Asking an AAC user to switch devices, potentially waiting months for improved communication, raises both logistical and ethical concerns. Kristensson et al. [25] have used a design engineering approach and presented a conceptual design for context-aware sentence retrieval intended for non-speaking individuals who have motor disabilities. In their study, they define both controllable and uncontrollable parameters for this design and examine the potential effectiveness of such systems through quantitative envelope analysis by varying the parameters. Hence, the design engineering method reveals insights about the feasibility of such systems without the immediate need to develop, introduce, and observe them for an extended duration. Similarly, Shen et al. [50] also used design engineering approaches to evaluate a multi-turn dialogue system for AAC using context-aware sentence generation by bag-of-keywords. The design engineering approach we propose in this paper shares a similar philosophy with their work.
2 Continual Learning
In a continual Learning setup, the learning system is exposed to a series of tasks, sequentially, over time. Each task is represented by a specific data distribution. The tasks can be related or completely different from each other. The new tasks can come from various domains depending on the application, such as new classes in image classification, new users in recommendation systems, or new environments in reinforcement learning. The model is expected to learn these new tasks while maintaining its performance on the old tasks. One of the key aspects of the continual learning setup is that the model typically does not have access to the data from previous tasks while learning a new one, due to memory constraints or privacy issues.
@@ 2 Continual Learning
When a model, previously trained on older tasks, is simply fine-tuned on new tasks, its accuracy on the old tasks rapidly deteriorates. It loses its proficiency in tasks it had previously mastered, effectively replacing the old knowledge with the new. This phenomenon is referred to as catastrophic forgetting. This problem occurs because conventional neural networks update their weights primarily based on the most recent data during training, a process that can result in the erasure of previously acquired patterns and knowledge. More specifically, gradient-based optimization algorithms prioritize minimizing the loss of the current training task, often disregarding previous task parameter settings. Although constraining parameter updates can alleviate this issue, it hampers the model’s ability to effectively learn new tasks. This creates a dilemma between stability (retaining knowledge of old tasks) and plasticity (learning new tasks), posing a challenge for continual learning [14].
@@
There are also other different learning paradigms, such as multi-task learning [69], transfer learning [60], meta learning [55] and online learning [46]. However, while most of these methods focus on plasticity, they usually fail to address catastrophic forgetting, leading to a dramatic decrease in the performance with old tasks [10].
3 Open-World Gesture Recognition
We start by characterizing the OWGR by three real-world cases in detail. We then formally formulate our problem.
@@ 3.1 Real-World Cases
We identify three general real-world cases (new context, new gesture, new user) for open-world gesture recognition.
@@
    New context is where each task represents different environmental/activity contexts, such as standing, walking, riding in a car, pushing stroller/cart, laying down etc. Therefore, we desire an OWGR system that is initialized on training data only from standing contexts to be able to incrementally learn on new contexts such as walking and riding in a car, while also preserving the knowledge from the previous standing contexts. These contexts can be further split into more detailed contexts. For example, standing includes standing with hand up to the chest level and standing with hand hanging, and walking includes walking with hand up to the chest level and walking with hand hanging. Our dataset includes this finer granularity of activity context. In practice, we need a task descriptor to inform the model of each task in the new context. In the real-world deployment, the description can be determined by the contextual information provided by other devices, such as location or an HMD (e.g., AR glasses).
@@
    New Gesture is where each task represents a subset of distinctive gestures to be recognized. Typically, a given application needs to recognize only a certain set of gestures. For example, one application only needs to recognize pinch and fist clench, a second application needs to recognize pinch and middle finger pinch, and a third application needs to recognize middle finger pinch and double pinch. We optimize the recognizer for each scenario. Another example is gesture customization, e.g. if the default gesture for the command “confirm” is a single pinch with the index finger, but a user wants to use the middle finger instead. The task descriptor here can be identified through the system setting (user-authorized gestures) or through the application identity (different applications with a distinct set of gestures).
  3. 3.
@@ 3.2 Problem Formulation
We formulated our problem more formally in this section. In task-incremental setting for open-world gesture recognition, the tasks (new context/gesture/user) come sequentially and the learner optimizes until convergence within each task [43, 13, 39]. This setting has the following assumptions:
@@
    The recognition model will have a multi-head configuration, as each task needs a separate output layer. Therefore, a task descriptor is also fed into the recognition model since the algorithm needs to know which head to use for that specific task. In practice, this task descriptor would be generated through various sensing modalities (e.g., activity recognition and location information [57]). For example, a classifier using egocentric video from a head-mounted display would estimate an activity context. We discuss the feasibility of this method in Section 7.3.
@@
$\frac{1}{\left|D_{train}\right|}\sum_{(x_{i},y_{i})\in D_{train}}\mathcal{L}(f(x_{i};\theta),y_{i})$ (1)
@@
4 Large-Scale Data Collection
We collected a large-scale dataset of inertial measurement unit (IMU) data consisting of 6 dimensions (3-axis accelerometer and 3-axis gyroscope). Figure 2 illustrates exemplary IMU signals collected.
(a) holding a cup or something else in the other hand (no)
(b) riding a car as a passenger (no)
(c) standing w hand up to the chest level (yes)
(d) stationary biking (yes)
Figure 2: Illustration of IMU signals for the four gestures, single pinch, double pinch, middle pinch, fist clench, under four different contexts. Yes and no in the bracket means if the participants is looking at the device or not. IMU signals corresponding to the same gesture can vary significantly under different contexts.

### v2

@@ Abstract
Providing users with accurate gestural interfaces, such as gesture recognition based on wrist-worn devices, is a key challenge in mixed reality. However, static machine learning processes in gesture recognition assume that training and test data come from the same underlying distribution. Unfortunately, in real-world applications involving gesture recognition, such as gesture recognition based on wrist-worn devices, the data distribution may change over time. We formulate this problem of adapting recognition models to new tasks, where new data patterns emerge, as open-world gesture recognition (OWGR). We propose the use of continual learning to enable machine learning models to be adaptive to new tasks without degrading performance on previously learned tasks. However, the process of exploring parameters for questions around when, and how, to train and deploy recognition models requires resource-intensive user studies may be impractical. To address this challenge, we propose a design engineering approach that enables offline analysis on a collected large-scale dataset by systematically examining various parameters and comparing different continual learning methods. Finally, we provide design guidelines to enhance the development of an open-world wrist-worn gesture recognition process.
@@ Introduction
Wrist-worn gesture recognition using inertial measurement unit (IMU) signals offers a convenient, always-on interface for various applications in mixed reality [21, 54, 42, 46]. Currently, most hand gesture recognition algorithms are optimized for closed-world settings, where training and test data come from the same underlying distribution [53]. However, in real-world applications [3, 43], the assumption that training and test data belong to the same underlying distribution no longer holds as: (1) new gesture data continuously arrives with changing characteristics; (2) gesture data may change over time; and (3) entirely new data patterns can emerge. We here call this problem open-world gesture recognition (OWGR). A static machine learning process, in which a recognition model is trained once and subsequently deployed, cannot effectively tackle this problem. What we seek is an adaptive machine learning process that can continuously train and deploy the model on newly emerging data, as illustrated in Figure Towards Open-World Gesture Recognition.
Various approaches can be used for adaptive machine learning. Simply retraining a recognition model with the entire joint dataset, past data and new data from new tasks, is sometimes infeasible due to limited computational power on embedded devices. Storing the entire past dataset is also challenging due to the limited memory of the device and privacy concerns. On the other hand, if a model trained from past data is naïvely fine-tuned on a new task, that model will dramatically decrease the recognition performance on the old tasks [47]. This is called catastrophic forgetting. Continual learning, sometimes called lifelong learning, methods are specifically designed to alleviate catastrophic forgetting by balancing the trade-off between plasticity (transferring knowledge from an old task to a new task) and stability (catastrophic forgetting).
We identify two real-world cases that fall under OWGR, new context and new user, as shown in Figure Towards Open-World Gesture Recognition. In all two cases, we wish to preserve performance on previously learned tasks, to avoid catastrophic forgetting, while at the same time learning the new task. Here, a new task may refer to a new context, or a new user. In Section 2, we present a key contribution in this paper—a detailed formulation of the open-world gesture recognition problem and its associated real-world scenarios, in terms of these two cases: (1) new context; and (2) new user.
@@
In a static machine learning process, we only need to log the data from the device, optimize the recognition model on the logged data to best fit the data once, and then deploy a trained model onto the device. However, in an adaptive machine learning process for OWGR, we additionally introduce additional stages as illustrated in Figure 1. In particular, this OWGR process includes two additional stages that require careful design considerations: (1) when to update the model; and (2) how to update the model. When refining a wrist-worn open-world gesture recognition system, developers face challenges due to the complexity of user studies and numerous parameters. For instance, evaluating new context requires recruiting participants to perform gestures in various real-world contexts, not just in a lab. Additionally, optimizing the OWGR process often needs multiple evaluations, making it difficult to execute and validate.
We therefore follow a prior approach [25], which proposes using a design engineering approach to assess a process that is challenging to evaluate via user studies. We have adopted this design engineering approach by translating it into a strategy that outlines a functional design of open-world gesture recognition and examines its conceptual design in which the key function of the process (such as Updating Policy) is translated into six potential function carriers (five different implementations of continual learning methods and one finetuning baseline). We then distinguish between controllable and uncontrollable parameters of these functions and assess the process’s effectiveness through quantitative envelope analysis. Moreover, given the lack of prior research that focuses specifically on OWGR, it is crucial to first apply the most advanced and generalizable continual learning methods to address this issue. Only after extensive exploration and evaluation of OWGR using other more generalizable and well-established continual learning methods are we in a position to best understand how to propose new algorithms to further enhance the system. Therefore, proposing a new algorithm is beyond the scope of this paper.
@@
Such a design engineering approach allows us to assess controllable and uncontrollable parameters offline, thus sidestepping the need for on-ground deployments and tests with real users in the first instance. To this end, we first collect a large-scale dataset using the IMU sensor from wrist-worn devices. To capture diversity in the aforementioned two real-world cases, the dataset contains 50 users performing four different gestures in 25 different contexts. We then construct a surrogate data model and a surrogate task model, identify their controllable and uncontrollable parameters, and subsequently perform an envelope analysis by varying these parameters, which allows us to assess emergent qualities of the system as a whole via simulation. This approach is then able to illustrate that using a continual learning method for open-world gesture recognition significantly improves gesture recognition accuracy and reduces catastrophic forgetting compared to a finetuning method. This forms our second contribution to the paper.
Designers using engineering methodologies to refine the OWGR process face limitations in data collection and computational resources. Envelope analysis, being computationally intensive, demands extensive simulations and parameter sweeps. Our study required over 40,000 GPU hours for dataset collection and offline evaluation of continual learning methods, highlighting the challenge for resource-constrained designers. To aid these designers, we provide essential design guidelines derived from our methodology. Our diverse dataset, covering various users and contexts, ensures the generalizability of these guidelines. Therefore, our third contribution is offering distilled guidelines for developing OWGR processes based on our engineering approach using envelope analysis.
In summary, this paper makes three contributions: (1) a detailed formulation of the open-world gesture recognition (OWGR) problem; (2) a design engineering approach for allowing open-world gesture recognition to be tractable in practice; and (3) design guidelines for developing OWGR processes.
@@ 1.1 Wrist-Worn Devices with Mixed Reality
To overcome this limitation, wrist-worn devices have been employed for gestural interaction with mixed reality devices [2, 13]. Gesture recognition is conducted using these wrist-worn devices, eliminating the need for HMD cameras to detect hands and gestures. Using wrist-worn devices facilitates gestural interactions even when hands are under the table or in a pocket, offering a seamless user experience. Moreover, wrist-worn devices can recognize gestures in a broader variety of ways compared to front-facing cameras, which primarily rely on computer vision techniques. There has been extensive research exploring hand gesture recognition using a wide range of wrist-worn sensing modalities, such as RGB cameras [64, 18, 65], infrared (IR) ranging [35], inertial measurement unit (IMU) [23, 17], acoustics [38, 27], electromyography (EMG) [47, 6, 36], electrical impedance tomography [71], pressure [12, 22], radar [29], stretch sensors [55], magnetic sensors [9, 39], and bio-capacitive effects [57]. IMU sensors are widely used for gesture recognition owing to their low cost and low power characteristics [73, 66].
1.2 Gesture Recognition
@@
As for gesture recognition techniques, early trajectory-based gesture recognition methods (e.g., dynamic time warping (DTW) [30] and hidden Markov models (HMM) [37]) can recognize simple gesture trajectories such as drawing a shape  [37]). However, these methods do not work well for more complex and fine-grained gestures such as pinching or making a fist. Recently, researchers and practitioners mainly use data-driven approaches by collecting a labeled gesture dataset. They then either train traditional models such as SVM, trees, (e.g., [16, 19]) or deep learning models when the dataset is large enough (e.g., [18, 68]). However, most prior work focuses on recognizing a pre-defined gesture set. There are very few works addressing open-world gesture recognition. Xu et al. proposed a few-shot learning framework for gesture authoring [66]. Shen et al. proposed a deep-learning model that can learn a new gesture with a synthetic dataset that is generated from a few data samples with a deep generative model [51, 49, 50]. Wang et al.  [62] identified a continual learning application for lifelong egocentric gesture recognition such that a VR system allows users to customize gestures incrementally. While these approaches are useful for models to learn new gestures, they are not appropriate for the two cases we identified, new context and new user. Moreover, Xu et al. [66] and Shen et al. [51] also fail to address the catastrophic forgetting problem when old gesture data are revisited.
1.3 Design Engineering
@@
Design engineering is a holistic method to create products and systems. This approach transitions from the initial stage of problem identification to tackling design-related aspects during the life cycle of a product or system, including production, upkeep, and eventual decommissioning. Validating certain systems, such as a a context-aware sentence retrieval system for AAC users, is extremely challenging [25] as such systems require tailored setups and extended use before benefits emerge. Asking an AAC user to switch devices, potentially waiting months for improved communication, raises both logistical and ethical concerns. Prior work [25] has used a design engineering approach to tackle this problem. This approach entails defining controllable and uncontrollable parameters for the design and examining the potential effectiveness of such systems through quantitative envelope analysis by varying the parameters through simulation [25]. Such a design engineering method reveals insights about the feasibility of such systems without the immediate need to develop, introduce, and observe systems for extended time periods. In addition, it allows examination of emergent qualities of systems that would be very difficult to study in empirical user studies, such as mechanisms explaining users strategies [26]. Similarly, Shen et al. [52] also used a design engineering approach to evaluate a multi-turn dialogue system for AAC using context-aware sentence generation by bag-of-keywords. The design engineering approach we propose in this paper shares a similar philosophy with this prior work work.
1.4 Continual Learning
@@
In a continual learning setup, the learning system is exposed to a sequence of tasks over time, each represented by a specific data distribution. The tasks may be related or completely different. The model is expected to learn new tasks while maintaining performance on previous ones, typically without access to data from earlier tasks due to memory constraints or privacy issues. When a model, previously trained on older tasks, is simply fine-tuned on new tasks, its accuracy on the old tasks rapidly deteriorates. It loses its proficiency in tasks it had previously mastered, effectively replacing the old knowledge with the new. This phenomenon is referred to as catastrophic forgetting. This problem occurs because conventional neural networks update their weights primarily based on the most recent data during training, a process that can result in the erasure of previously acquired patterns and knowledge. More specifically, gradient-based optimization algorithms prioritize minimizing the loss of the current training task, often disregarding previous task parameter settings [15].
@@
There are also other different learning paradigms, such as multi-task learning [72], transfer learning [63], meta learning [58] and online learning [48]. However, while most of these methods focus on plasticity, they usually fail to address catastrophic forgetting, leading to a dramatic decrease in the performance with old tasks [11]. Being able to avoid catastrophic forgetting is critical in OWGR as it is not solely about quickly learning new data - a challenge that can be addressed by few-shot learning [56] or domain adaptation [34] - but more importantly, the core challenge of OWGR is to retain old data. More specifically, domain adaptation focuses on adapting to new domains without necessarily retaining performance on the original domain, which does not directly address catastrophic forgetting. Few-shot learning deals with learning new tasks or adapting to new data quickly from few examples, but it does not inherently address learning multiple tasks sequentially without forgetting. These two methods are specifically designed to optimize plasticity (transferring knowledge from an old task to a new task), without considering stability (catastrophic forgetting). Therefore, these learning paradigms are not well-suited to the OWGR problem. Hence, we do not directly compare continual learning methods with other learning paradigms in this paper.
2 Open-World Gesture Recognition (OWGR)
@@ 2.1 Real-World Cases
    New context is where each task represents different environmental/activity contexts, such as standing, walking, riding in a car, pushing stroller/cart, laying down etc. Therefore, we desire an OWGR system that is initialized on training data only from standing contexts to be able to incrementally learn on new contexts such as walking and riding in a car, while also preserving the knowledge from the previous standing contexts. These contexts can be further split into more detailed contexts. For example, standing includes standing with hand up to the chest level and standing with hand hanging, and walking includes walking with hand up to the chest level and walking with hand hanging. Our dataset includes this finer granularity of activity context. In practice, we need a task descriptor to inform the model of each task in the new context. In the real-world deployment, the description can be determined by the contextual information provided by other devices, such as location or an HMD.
@@ 2.2 Problem Formulation
We formulate the OWGR problem more formally in this section. In a task-incremental setting for open-world gesture recognition, the tasks (new context/gesture/user) arrive sequentially and the learner optimizes until convergence within each task [45, 14, 41]. This setting has the following assumptions:
@@
    The recognition model will have a multi-head configuration, as each task needs a separate output layer. Therefore, a task descriptor is also fed into the recognition model since the algorithm needs to know which head to use for that specific task. In practice, this task descriptor would be generated through various sensing modalities (e.g., activity recognition and location information [60]). For example, a classifier using egocentric video from a head-mounted display would estimate an activity context. We discuss the feasibility of this method in Section 5.2.1 and Section 5.5. Despite a multi-head configuration may not be necessary for the new user case, the multi-head approach simplifies the comparative analysis of different continual learning methods — allowing for consistent evaluation with minimal adjustments.
@@
$\frac{1}{\left|D_{train}\right|}\sum_{(x_{i},y_{i})\in D_{train}}\mathcal{L}(f(x_{i};\theta),y_{i}),$ (1)
@@
2.3 Large-Scale Data Collection
We collected a large-scale dataset of inertial measurement unit (IMU) data consisting of 6 dimensions (3-axis accelerometer and 3-axis gyroscope).  [^1] Figure 1 in the Appendix illustrates exemplary IMU signals collected.